### v1

# Photodissociation spectra of single trapped CaOH molecular ions

Abstract
Molecular ions that are generated by chemical reactions with trapped atomic ions can serve as an accessible and successful testbed for developing molecular quantum technologies. On the other hand, they are also a hindrance to scaling up quantum computers based on atomic ions as unavoidable reactions with background gas destroy the information carriers. Here, we investigate the single-photon and two-photon dissociation processes of single CaOH molecular ions co-trapped in Ca ion crystals using a femtosecond laser system. We report the photodissociation cross section spectra of CaOH for single-photon processes at nm and for two-photon processes at nm. This result can serve as a basis for dissociation-based spectroscopy for studying the internal structure of CaOH. The result also gives a prescription for recycling Ca ions in large-scale trapped Ca quantum experiments from undesired CaOH ions formed in the presence of background water vapor.
I Introduction
Molecules possess various internal degrees of freedom that are both chemically and physically intriguing. Molecular ions in radio-frequency traps enable precision molecular spectroscopy on the single molecule scale and are thus a promising platform for applications of molecules in quantum technologiesDeMille (2002); Lin et al. (2020) and the exploration of fundamental physicsRoussy et al. (2023); Alighanbari et al. (2020). For molecular ions, direct laser cooling via cycling transitions proves to be challenging and has not yet been realized experimentallyNguyen et al. (2011); Ivanov et al. (2020); Wójcik, Hudson, and Krylov (2022). While the translational motion of trapped molecular ions can be sympathetically cooled by laser cooling co-trapped atomic ionsMølhave and Drewsen (2000); Rugango et al. (2015); Wan et al. (2015), the absence of suitable cycling transitions poses a challenge in investigating the internal structure of molecular ions. Most spectroscopic experiments with trapped molecular ions rely on quantum logic methods utilizing co-trapped atomic qubitsWolf et al. (2016); Chou et al. (2017); Sinhal et al. (2020); Chou et al. (2020) or destructive detection methods based on, e.g., photodissociation channelsKoelemeij et al. (2007); Karr, Douillet, and Hilico (2012); Seck et al. (2014); Ni et al. (2014); Khanyile, Shu, and Brown (2015); Rugango et al. (2016).
Various molecular ions can be generated from trapped atomic ions colliding with background gas molecules in a vacuum chamber. Photodissociation of molecular ions generated in this way leads to the recovery of atomic ions. This can be observed from the increase in the overall atomic ion fluorescenceOkada et al. (2003); Bertelsen, Jørgensen, and Drewsen (2006), which provides a way to study the rovibrational structure of such molecular ions.
For trapped ion experiments based on Ca, which is one of the most well-studied and widely implemented ion species for quantum information processingSchindler et al. (2013), molecular ions generated in chemical reactions with background gas are predominantly CaH, CaO, and CaOHHansen et al. (2012). While extensive theoretical and experimental research has been conducted on the two diatomic speciesAbe et al. (2012); Rugango et al. (2016); Chou et al. (2017); VanGundy, Bartlett, and Heaven (2018); Chou et al. (2020); Qi, Reed, and Brown (2023), there exists little spectroscopic study concerning the triatomic CaOH molecular ion.
As illustrated in Fig. 1.(a), CaOH can form in a trapped Ca ion experiment in the presence of water vapor in the vacuum chamber by the reactionOkada et al. (2003)
While CaH and CaO can be photodissociated with the 397 nm and 375 nm lasers used in trapped Ca experimentsWu et al. (2021), the photodissociation of CaOH requires a shorter wavelength for exciting the molecule to its unbound first electronic excited state. This photodissociation channel was studied by Okada et al. in 2006 using a combination of optical filters and a broadband UV lamp, where they estimated the dissociation threshold to be Okada et al. (2006). However, it was later reported by Hansen et al. that Ca can also be recovered from CaOH by applying 272 nm lightHansen et al. (2012). In this paper, we report on a measurement of the photodissociation spectrum of CaOH, where the wavelength-dependence of the photodissociation cross section was measured by excitation with a femtosecond optical parametric amplifier (OPA) system. With the same ultrafast laser system, we are also able to investigate two-photon dissociation of CaOH and the measured two-photon dissociation spectrum is reported. The result implies the possibility of recycling Ca from CaOH with ablation lasers at 515 nm, which are commonly used in trapped Ca quantum experiments.
II Experimental setup
As illustrated in Fig. 2, our experiment is based on a linear Paul trapPogorelov et al. (2021) in ultrahigh vacuum (UHV). A gate valve separates the chamber into two parts: the main experiment chamber and the molecular gas chamber. The main experiment chamber containing the ion trap is connected to a non-evaporable getter (NEG) pump and a NEG-ion combination pump to maintain a base pressure mbar. Ca ions in the ion trap originate from an integrated Ca target mounted near the trap. By ablating the target with 532 nm laser pulses, neutral Ca atoms are introduced to the ion trap where they are photoionized in a two-step process using 375 nm and 422 nm lasersPogorelov et al. (2021). The generated Ca ions are then captured by the ion trap and Doppler-cooled by driving their 4S 4P cycling transition at 397 nm, as shown in the inset of Fig. 2. Additional lasers at 866 nm and 854 nm are applied to repump the ions that decay to the 3D and 3D states, respectively, back to the cyclic cooling manifold.
To the right of the main chamber, the molecular gas chamber is used to accumulate water vapor for the generation of CaOH molecular ions. When the gate valve is closed, this section is not pumped by the vacuum pumps attached to the main chamber and the water vapor pressure inside increases over time due to outgassing from the chamber wall. To increase the partial pressure of water further, water vapor that is outgassing from an unbaked stainless steel hose is introduced to the molecular gas chamber via a leak valve.
III Molecule Generation and identification
To generate CaOH for the photodissociation measurement, water vapor is introduced to the main chamber by opening the gate valve. In our setup, the partial pressure of water vapor in the system cannot be directly measured. Instead, the increase in water vapor pressure in the main chamber is inferred from the change in total pressure estimated from the current of the ion pump. While waiting for the generation of CaOH, the gate valve is opened every 5 minutes for 1 second, which results in the total pressure estimated by the ion pump increasing from mbar to mbar.
As investigated by Okada et al. in 2003, the chemical reaction rate between HO and the trapped Ca ions increases when the ions are pumped into the 3D state by switching off the 866 nm repumping light during laser coolingOkada et al. (2003). Therefore, the 866 nm laser is periodically switched off for 1 second every 2 seconds in the molecule generation phase. The generation rate of one CaOH molecular ion is typically 10 to 15 minutes for the ion crystals of Ca ions investigated in this experiment.
The above molecule generation process was found to produce predominantly CaOH by mass spectrometry using the method presented by Drewsen et al. Drewsen et al. (2004) (for details, refer to Appendix D). In this method, the unknown molecular mass can be inferred from the frequency of the axial center-of-mass (COM) motion of a 2-ion crystal of one Ca ion and one molecular ionKielpinski et al. (2000); Drewsen et al. (2004) as
where is the ratio between the atomic and molecular mass and is the COM motional frequency of a Ca ion crystal. These frequencies were measured by applying an oscillating voltage to one of the endcap electrodes. When the frequency of this “tickling” signal is resonant with that of the COM motion of the ion crystal, it excites the corresponding oscillation of the crystal. This increases the motional temperature of the ion crystal and thus reduces the scattered Doppler cooling light. The COM motional frequency is thus determined from the ion crystal’s fluorescence as a function of the tickling frequency. The masses of the generated molecular ions are measured with a precision of m amu and the results are shown in Fig. 3. All 65 measurements have masses consistent with CaOH, from which we conclude that the molecular ions generated in our procedure are predominantly this species.
IV Photodissociation
After verifying the generation of a single CaOH, the measurement of the photodissociation spectrum of CaOH is performed by applying femtosecond pulse trains produced by an OPA. The spatial beam profile and position of the OPA light are measured by a beam profiler mounted on a rail. To assist the alignment of the OPA light, an additional 397 nm Doppler cooling beam is applied co-propagating with the OPA light, which is aligned to the center of the ion trap by maximizing the 397 nm light scattered by the ion crystal. The OPA light is overlapped with this axial Doppler cooling beam at the position of the ions. We estimate the alignment precision of this procedure to be 10 µm, limited by the resolution of the beam profiler. The radius of the OPA light at the ion crystal was also estimated by the beam profiler to be 150(10) µm.
We assume that the dissociation of a CaOH molecule during illumination is memoryless due to the lack of intermediate electronic states. Thus, the dissociation probability after illumination time can be described by the cumulative distribution function (CDF) of an exponential distribution,
characterized by the mean dissociation time , where is the wavelength of the OPA light. The effective photodissociation cross section can be related to the mean dissociation time by (see Appendix A.1)
Here, is the photon energy, is the Planck’s constant, is the speed of light, and is the average intensity of the OPA light at the ion position with average power and beam radius .
The experimental procedure is initialized by loading a one-dimensional ion chain with Ca ions into the ion trap for the generation of a single CaOH (see Fig. 10). While Ca ions can be detected on an electron-multipling CCD (EMCCD) camera, the generated CaOH ions do not scatter light and are recognized as dark defects in the ion chain. The number of dark ions is identified by comparing the ion crystal image with pre-calculated ion crystal configuration templates with different numbers of dark and bright ions (see Appendix C). After the appearance of a single molecular ion, the OPA light is applied to illuminate the molecular ion for up to 300 s. If the OPA light successfully recovers the original Ca from the CaOH ion, as shown in Fig. 4, the illumination time for achieving photodissociation will be recorded. Otherwise, if the photodissociation is not successful or the dark ion is lost during the process, the dissociation trial is marked as unsuccessful within the accumulated illumination time. In either case, utilizing Bayesian analysis, the -th trial provides information to update the prior distribution of the mean dissociation time by multiplying it with a likelihood function
based on Eq. (2) where denotes the total illumination time. Starting from a uniform prior distribution in the range of s, the posterior distribution after trials,
reveals an estimation of . To ensure convergence of the posterior distribution, approximately 20 trials were performed at each wavelength.
As shown in Eq. (3), linear dependence of the dissociation rate on the average power of the OPA light is expected. This is verified in the measurement of the dissociation rate at different at nm. Fitting the results to a power law function, as shown in Fig. 5.(a), reveals an exponent of . Thus, to better estimate the mean dissociation time, the average power of the beam is adjusted between 6 and 60 µW in order for at least half of the dissociation trials to be successful within the maximum illumination time. The repetition rate of the OPA is fixed at 1 kHz. The photodissociation cross sections at wavelengths ranging from 245 to 275 nm are then measured and shown in Fig. 5.(b).
V Two-photon dissociation
In addition to the measurement of single-photon dissociation, we also investigated the photodissociation of CaOH at wavelengths around 500 nm. To achieve a dissociation rate that is comparable to the single-photon measurements, the repetition rate of the OPA is increased to 10 kHz and the light intensity is increased by reducing the beam radius to 70(10) µm and increasing the average power to a few milliwatts. At nm, the dependence of the dissociation rate on the average intensity of the OPA light on the ions is investigated, as illustrated in Fig. 6.(a). Fitting the data to a power law function gives an exponent of , which is consistent with two-photon processes.
Thus, the photodissociation of CaOH at wavelengths around 500 nm is assumed to be dominated by a two-photon dissociation channel via its unbound first electronic excited state. The two-photon dissociation cross section is related to the dissociation rate by (see Appendix A.2)
which gives
in units of GM (1 GM = ), where a Gaussian temporal pulse profile is assumed and is the full width at half maximum (FWHM) pulse duration. The peak intensity is related to the average intensity by . The pulse duration at each wavelength is assumed to be Fourier-transform-limited and is thus estimated from the spectral bandwidth of the OPA light (see Appendix B.2). The two-photon dissociation cross section spectrum at nm is then measured and shown in Fig. 6.(b).
VI Discussion
In summary, we measured the photodissociation spectrum of single trapped CaOH molecular ions sympathetically cooled in Ca ion crystals at nm and nm. The photodissociation processes at wavelengths in these two ranges are verified to originate predominantly from a single-photon and a two-photon dissociation channel, respectively. While the one-photon dissociation cross section started to drop off around nm, a similar decrease was observed in the two-photon dissociation spectrum around nm. This implies that both processes involve the excitation of the molecular ion to its unbound first electronic excited state which leads to its dissociation. Compared to the previous workOkada et al. (2006) on single-photon dissociation of CaOH by Okada et al. in 2006, we report a more precise measurement of the dissociation cross sections at various wavelengths which is consistent with the previous estimation of at nm. The measured single-photon dissociation spectrum is also consistent with the photodissociation of CaOH at 272 nm observed by Hansen et al.Hansen et al. (2012).
As for the two-photon dissociation of CaOH around nm, the dissociation process is found to be significantly less efficient compared to the single-photon dissociation process, as expected since higher peak intensity is typically required for higher-order processes. Yet femtosecond lasers at such wavelengths have proven to be alternatives to continuous wave lasers with nm for photodissociation of the molecular ion. The question arises as to whether nanosecond lasers that are used for ablation loading in trapped atomic ion experiments at nmPogorelov et al. (2021) are able to dissociate CaOH. Achieving a mean photodissociation time of 1 minute for a single CaOH ion is estimated to require an average power of W, considering a laser with kHz, ns, and the light focused onto the ions with a beam waist of µm.
The reported measurement is implemented on single CaOH molecular ions in small trapped ion crystals, which enables the relatively precise evaluation of the photodissociation cross section due to the repeatable positioning of the molecule. The scale of the system is also crucial for investigating a two-photon dissociation process that requires tightly focused light to achieve sufficient intensity. The reported photodissociation spectrum can be a basis for studying the rovibrational structure of CaOH via dissociation-based spectroscopy. Designed to investigate the interaction between femtosecond laser pulses and a linear ion crystal with a single CaOH, this experiment represents a precursor for performing quantum logic spectroscopy with polyatomic molecular ions Schindler (2019).
Acknowledgements
This research was funded by ERC-2020-STG 948893, ESQ Discovery project SDEF: State-dependent force spectroscopy for trapped ions, and FWF 1000 Ideas project TAI-798. The authors would like to acknowledge AQT for support with the ion trap and vacuum pumps, in particular Daniel Nigg and Georg Jacob; Markus Teller for assistance with vacuum work; electronics workshop technicians Kilian Prokop and Wolfgang Kuen for lab infrastructure support; Milan Oncak for discussions on quantum chemistry calculations; Christian Marciniak for general discussions; Marco Valentini for Ca lasers implementation; and the Quantum Optics & Spectroscopy Group and associated ion trapping groups at the Universität Innsbruck for general assistance.
Author Contributions Z.W., S.W., B.F., and P.S. designed the experiment. Z.W., S.W., and B.F. carried out the dissociation experiment and analyzed the data. M.I.M., E.M., and B.F. performed the characterization of the OPA system. Z.W., S.W., and B.F. contributed to the manuscript. All authors contributed to the experimental setup and reviewed the manuscript. B.F. and P.S. supervised the project.
Conflict of Interest
The authors have no conflicts of interest to disclose.
Data and Code Availability
The data and analysis code that support the findings of this study are openly available at https://github.com/quantum-molecules/dissociation_analysis.
Appendix A Photodissociation model
A.1 Single-photon dissociation
The photodissociation process considered in this paper is treated as a consequence of a photon absorption event which excites the molecular ion to its unbound first electronic excited state. We thus define the effective photodissociation cross section in terms of the absorption cross section and the quantum yield . We assume all of the photon absorption events lead to dissociation of the molecular ion, i.e., , and the molecular rovibrational structure is not considered for simplicity.
The absorption cross section of a single molecule is related to the absorption coefficient of molecular gas by
where denotes the molecular number density. For light propagating through the absorptive molecular gas, the change in light intensity is described by the differential equation
Hereby we consider photons absorbed by molecules in a small fraction of the molecular gas with volume and the photon absorption event in each molecule that leads to the excitation of the molecule is considered as independent. Regarding a train of ultrafast laser pulses separated in time by , the mean number of excitations induced by each laser pulse can be expressed as
where is the photon energy and is the number of molecules.
As the measured dissociation time is at a larger time scale, the mean number of excitations per unit time is given as
where and . Considering a single molecule () that can only be excited once, the probability of the molecule being excited per unit time is
where is then the exponential distribution CDF,
Therefore, the mean dissociation time can be related to the absorption cross section and thus the effective dissociation cross section by
A.2 Two-photon dissociation
Similarly, we define the effective two-photon dissociation cross section to be , related to the two-photon absorption cross section and the quantum yield . Considering two-photon absorption in a molecular gas, the change in light intensity is described by
characterized by the two-photon absorption coefficient , where is the two-photon absorption cross section in units of GM. Compared to single-photon dissociation, a two-photon dissociation event requires the absorption of two photons to excite the molecule to its unbound excited state. Similar to Eq. (10), for a train of ultrafast laser pulses separated in time by , the mean number of excitations induced by each laser pulse can be expressed as
Assuming the pulses have a Gaussian temporal profile, i.e., , the mean number of excitations per unit time can be derived as
where is the FWHM pulse duration. Similar to the single-photon case in Eq. (14), the mean dissociation time of a single molecule is thus related to the two-photon dissociation cross section by
A.3 Memory of the photodissociation process
To verify the memorylessness of the photodissociation process, the illumination for photodissociation is divided into several time intervals with various fixed durations. From this data, we can estimate the probability of dissociation for a fixed illumination duration as well as the cumulative time until dissociation. If the photodissociation process has no memory effect, the probability of the molecule being dissociated after being illuminated for a fixed time will be described by the exponential distribution CDF as in Eq. (2).
We can therefore also estimate the mean dissociation time using the estimated dissociation probability for fixed times. These probabilities are then fitted to the exponential distribution CDF as indicated in Fig. 7, e.g., for nm and nm. We estimate the photodissociation cross sections via this analysis for all data and compare them to the cumulative dissociation time, as shown in Fig. 8. We find that the cross sections obtained from both methods agree within the uncertainties which indicates that the single- and two-photon dissociation processes are memoryless.
Appendix B Experimental details
B.1 Ca Lasers and trap parameters
During the photodissociation measurement, the cooling laser at 397 nm and the repumpers at 866 nm and 854 nm are constantly applied for Doppler cooling the ion crystal. The upper limit of the power applied for each laser is 0.6 mW (397 nm), 0.4 mW (866 nm), and 0.3 mW (854 nm). The beam radii of the lasers at the ion position are larger than µm. To provide axial confinement of the ion crystal, the voltage on the endcap electrodes of the ion trap is set to 205 V. This leads to an axial COM motional frequency of kHz for all the experiments in this paper, based on the measurement discussed in Appendix D.
B.2 Autocorrelation and spectra of OPA
The femtosecond OPA system used in this work is an ORPHEUS-HP OPA seeded and pumped by a Carbide CB5 amplifier from Light Conversion. As knowledge of the pulse duration is required to obtain the two-photon dissociation cross section, we characterized it using a combination of measurements of the spectra and autocorrelation measurements. While the pulse duration at nm could not be directly measured with our current setup, the pulses produced by the OPA are guaranteed to be transform-limited by the manufacturer. To verify this, the pulse duration and the spectrum of the OPA at nm were measured and the time-bandwidth product was inferred, as shown in Fig. 9.
To measure the pulse duration, we used an autocorrelator setup that splits the light from the OPA into two paths, where one path has an adjustable time delay. Overlapping both branches on a phase-matched BBO crystal induces type I second-harmonic generation. Assuming a Gaussian temporal profile, the FWHM of the SHG signal was measured to be fs, corresponding to a pulse duration of fs.
The spectra of the OPA light were measured with an Ocean Insight Ocean HR spectrometer. Fitting the spectrum at nm to a Gaussian distribution, the FWHM spectral width was determined to be THz. This gives a time-bandwidth product of , which is in agreement with the time-bandwidth product of a transform-limited Gaussian pulse, i.e., .
Thus, we assume the pulses are transform-limited at nm in the two-photon dissociation experiments. The measured center wavelengths, their corresponding spectral bandwidths, and the inferred pulse durations are reported in Table 1.
| (nm) | (THz) | (fs) |
| 501.63(1) | 3.83(1) | 96.8(3) |
| 510.16(1) | 4.20(2) | 91.2(4) |
| 520.526(6) | 4.096(7) | 97.4(2) |
| 527.339(8) | 3.47(1) | 118.0(3) |
| 530.778(5) | 2.963(6) | 140.0(3) |
| 536.025(7) | 3.489(9) | 121.3(3) |
| 540.597(8) | 3.06(1) | 140.6(4) |
B.3 Error estimation
In addition to the statistical uncertainties in the measured parameters of OPA pulse duration, the following sources of uncertainty are considered. The uncertainties of the mean dissociation times are given by Bayesian analysis, which are significantly larger than the uncertainties of the measured dissociation times. The uncertainty of the average intensity of the OPA light at the ion crystals derives from the uncertainty in average power (µW in single-photon measurements, mW in two-photon measurements), the uncertainty of beam positioning of (µm), and the uncertainty of the beam radius. The beam radius is estimated by the beam profiler with a precision of µm at a position corresponding to the real ion crystal position in the beam path. The deviation from the real ion crystal position is mm and results in a possible systematic error in the measured beam radius at the ion crystal, which is µm. The total uncertainty of the beam radius is thus considered to be µm.
B.4 Confidence
The uncertainties presented in this paper are based on a 68% confidence level.
B.5 Photodissociation experiment process tree
The photodissociation experiments in this paper are carried out with an automated script according to the process tree illustrated in Figure. 10.
Appendix C Ion crystal configuration identification
The photodissociation experiment requires real-time evaluation of the ion crystal configuration such that the number of fluorescing (bright) Ca and non-fluorescing (dark) molecular ions can be monitored. Knowledge of the ion crystal configuration allows a script to run automatic measurements and take actions such as (re)loading the trap or emptying it if more than one dark ion is created. The ion crystal configuration is identified using only image data of the ions taken by an imaging system with magnification using an Andor iXon+ 885 camera with a pixel size of 8 µm. For this process, a region of interest (ROI) in the image centered around the ions is converted into two normalized data sets: the column sum, where the ROI is summed along the columns of the image array; and the row sum, where the summation is done along the rows (see Fig. 11). All analysis to identify the ion crystal configuration is based on these two data sets.
The identification of the ion crystal configuration is performed in real-time and runs continuously throughout the measurements. An image of the ions is taken with an exposure time of 0.2 s, from which the ion crystal configuration is identified. The identification process takes s, then the next image is taken and analyzed. Three algorithms are used to determine the ion crystal configuration: (i) check for the presence of bright ions; (ii) counting of bright ions; and (iii) determination of the ion crystal configuration via a template fitting method.
The first method checks for the presence or absence of bright ions in the trap. Here, a moving average is applied to the row sum of the image data before applying a discrete Fourier transform (DFT). If the trap is empty, the moving average is mainly flat and the amplitude of the longest spatial wavelength in the DFT is small; if there are bright ions in the trap, the moving average exhibits a broad peak, and thus the amplitude of the longest spatial wavelength in the DFT is larger. A threshold is then set to differentiate between an empty or a non-empty trap.
The number of bright ions is counted from the row sum via a peak finder algorithm, which can determine the number and positions of bright ions without requiring a fluorescence-based threshold. This algorithm identifies the largest amplitude difference in a set of sorted local maxima (peaks) of row sum and uses this to classify peaks into two distinct groups: large-amplitude peaks originating from bright ions and small-amplitude peaks due to noise. Knowledge of the positions of the bright ion peaks can be utilized to identify the presence of dark ions for, e.g., determining the trap center position, as it requires an ion crystal of solely bright ions.
In order to count dark ions, a template fitting procedure was performed to identify the ion crystal configuration. In this process, template sets for different ion numbers and configurations are compared to the column sum data set. Template sets are generated by the following scheme. The desired ion configuration in a template for ions is given by the vector . The value of is assigned to for a bright Ca ion and for a dark molecular ion, for , e.g., the ion crystal configuration template which matches the row sum shown in Fig. 11 is . Then the sum of Gaussian peaks
is discretized by performing the set of integrals
Here denotes the numerically calculated positionsJames (1998) of the ions with respect to the calibrated trap center position , is the waist of the Gaussian peaks (taken from the fit presented in Fig. 11), and is an offset of the trap center position. The start and end positions, and , of the -th binning interval are chosen such that the binning and length of the template match those of the measured column sum data set. is the binned template of ions for the configuration . To find the particular ion crystal configuration, a large set of templates is created for all reasonable ion numbers and all possible corresponding ion configurations. All templates are compared with the column sum data set using the comparison function
where denotes the column sum data set and is its size. The best-fitting template, given by the minimum outcome of , then determines the ion crystal configuration. This method reliably counts the total, bright, and dark ion numbers and thus enables automation of the photodissociation and mass spectrometry measurements. Since this method is vulnerable to drifts of the trap center position, every particular template is scanned over column sum by varying the offset . This allows tracking of the trap center position and automated recalibration when the drift exceeds a given threshold.
Appendix D Mass spectrometry
A radio frequency (RF) “tickling” signal with an amplitude of 30 mV modulates the DC voltage on the endcap electrodes, which defines the axial confinement, and excites the COM motion of the ions which peaks when on resonance with the COM motional frequency. Due to Doppler broadening, the ions fluoresce less when their motional state is excitedWineland, Drullinger, and Walls (1978). The COM motional frequency can then be determined by scanning the RF tickling frequency to minimize the fluorescence.
In crystals with mixed ion species, the COM motional frequency depends not only on the different ion masses but also on the specific ion crystal configurationHome (2013). For instance, the two ion configurations (Ca, Ca, CaOH) and (Ca, CaOH, Ca) have a different COM motional frequency, with a deviation of 1.7 kHz in our system. When applying an RF signal which is close to the COM motional frequency of the current ion configuration, the ions reconfigure their respective positions due to the induced motion. After the reconfiguration, the RF signal is no longer resonant with the COM motion which makes it difficult to perform a precise measurement of the resonance frequency. Thus, we perform measurements with crystals consisting of exactly two trapped ions, as there exists only a single COM motional frequency due to symmetry. However, the dissociation measurements are performed with larger ion crystals to speed up the CaOH generation process. In order to prevent ion loss during the tickling scan, a Doppler cooling cycle of about is applied to the ions with the RF signal being turned off after each tickling measurement in the scan. For every single Ca and molecular ion pair, we performed and summed up to five measurement runs.
In order to determine the COM motional frequency from the images taken with the EMCCD camera, the measured photon counts at each pixel are summed in a small ROI around the ions. This ROI is wide along the y-axis of the images, where is the waist of a Gaussian fitted to the row sum dataset, as shown in Fig. 11. Along the x-axis, this small ROI has a length which is 5 times that of a two ion crystal with an endcap electrode voltage of 205 V, i.e., 8.5 µm. The total photon counts integrated over the ROI as a function of the RF tickling frequency are normalized and the COM motional frequency is determined by fitting a Gaussian function to the data, as shown in Fig. 12.
The fitted COM motional frequencies from the 65 Ca and molecular ion pairs and the corresponding reference COM motional frequencies of Ca gave masses for the molecular ions ranging from amu. This confirms that all measured 65 molecular ions were CaOH.
In order to estimate the statistical uncertainty of the measured COM motional frequency of trapped atomic Ca ions, we evaluated 34 data sets taken over the course of hours on two different days. For a voltage of 205 V applied to the endcap electrodes, a COM motional frequency of 526.8 kHz with a standard deviation of 200 Hz was determined. Since the errors in the fitted COM motional frequency of all 65 processed datasets of Ca and molecular ion pairs and all reference datasets are all smaller than 200 Hz, this value is utilized to obtain an upper limit for the uncertainty of COM motional frequencies. This results in an estimated uncertainty in the derived molecular ion masses of 0.1 amu.

### v2

@@
-Photodissociation spectra of single trapped CaOH molecular ions
+Photodissociation spectra of single trapped CaOH+ molecular ions
 Abstract
-Molecular ions that are generated by chemical reactions with trapped atomic ions can serve as an accessible and successful testbed for developing molecular quantum technologies.
+Molecular ions that are generated by chemical reactions with trapped atomic ions can serve as an accessible testbed for developing molecular quantum technologies.
 On the other hand, they are also a hindrance to scaling up quantum computers based on atomic ions as unavoidable reactions with background gas destroy the information carriers.
-Here, we investigate the single-photon and two-photon dissociation processes of single CaOH molecular ions co-trapped in Ca ion crystals using a femtosecond laser system.
-We report the photodissociation cross section spectra of CaOH for single-photon processes at nm and for two-photon processes at nm.
-This result can serve as a basis for dissociation-based spectroscopy for studying the internal structure of CaOH.
-The result also gives a prescription for recycling Ca ions in large-scale trapped Ca quantum experiments from undesired CaOH ions formed in the presence of background water vapor.
+Here, we investigate the single- and two-photon dissociation processes of single CaOH+ molecular ions co-trapped in Ca+ ion crystals using a femtosecond laser system.
+We report the photodissociation cross section spectra of CaOH+ for single-photon processes at nm and for two-photon processes at nm.
+Measurements are interpreted with quantum-chemical calculations, which predict the photodissociation threshold for at 265 nm.
+This result can serve as a basis for dissociation-based spectroscopy for studying the internal structure of CaOH+.
+The result also gives a prescription for recycling Ca+ ions in large-scale trapped Ca+ quantum experiments from undesired CaOH+ ions formed in the presence of background water vapor.
 I Introduction
@@
 (2003); Bertelsen, Jørgensen, and Drewsen (2006), which provides a way to study the rovibrational structure of such molecular ions.
-For trapped ion experiments based on Ca, which is one of the most well-studied and widely implemented ion species for quantum information processingSchindler et al.
-(2013), molecular ions generated in chemical reactions with background gas are predominantly CaH, CaO, and CaOHHansen et al.
+For trapped ion experiments based on 40Ca+, which is one of the most well-studied and widely implemented ion species for quantum information processingSchindler et al.
+(2013), molecular ions generated in chemical reactions with background gas are predominantly CaH+, CaO+, and CaOH+Hansen et al.
 (2012).
@@
 (2017); VanGundy, Bartlett, and Heaven (2018); Chou et al.
-(2020); Qi, Reed, and Brown (2023), there exists little spectroscopic study concerning the triatomic CaOH molecular ion.
-As illustrated in Fig. 1.(a), CaOH can form in a trapped Ca ion experiment in the presence of water vapor in the vacuum chamber by the reactionOkada et al. (2003)
-While CaH and CaO can be photodissociated with the 397 nm and 375 nm lasers used in trapped Ca experimentsWu et al.
-(2021), the photodissociation of CaOH requires a shorter wavelength for exciting the molecule to its unbound first electronic excited state.
+(2020); Qi, Reed, and Brown (2023), there are few spectroscopic studies concerning the triatomic CaOH+ molecular ion.
+As illustrated in Fig. 1.(a), CaOH+ can form in a trapped Ca+ ion experiment in the presence of water vapor in the vacuum chamber by the reactionOkada et al. (2003)
+While CaH+ and CaO+ can be photodissociated with the 397 nm and 375 nm lasers used in trapped Ca+ experimentsWu et al.
+(2021), the photodissociation of CaOH+ requires a shorter wavelength for exciting the molecule to its unbound first electronic excited state.
 This photodissociation channel was studied by Okada et al.
@@
 However, it was later reported by Hansen et al.
-that Ca can also be recovered from CaOH by applying 272 nm lightHansen et al.
+that Ca+ can also be recovered from CaOH+ by applying 272 nm lightHansen et al.
 (2012).
-In this paper, we report on a measurement of the photodissociation spectrum of CaOH, where the wavelength-dependence of the photodissociation cross section was measured by excitation with a femtosecond optical parametric amplifier (OPA) system.
-With the same ultrafast laser system, we are also able to investigate two-photon dissociation of CaOH and the measured two-photon dissociation spectrum is reported.
-The result implies the possibility of recycling Ca from CaOH with ablation lasers at 515 nm, which are commonly used in trapped Ca quantum experiments.
+In this paper, we report on a measurement of the photodissociation spectrum of CaOH+, where the wavelength-dependence of the photodissociation cross section was measured by excitation with a femtosecond optical parametric amplifier (OPA) system.
+With the same ultrafast laser system, we are also able to investigate two-photon dissociation of CaOH+ and the measured two-photon dissociation spectrum is reported.
+The result implies the possibility of recycling Ca+ from CaOH+ with ablation lasers at 515 nm, which are commonly used in trapped Ca+ quantum experiments.
+We employ quantum-chemical calculations to reproduce the measured photodissociation energy and predict the spectrum of excited electronic states of CaOH+.
 II Experimental setup
@@
 The main experiment chamber containing the ion trap is connected to a non-evaporable getter (NEG) pump and a NEG-ion combination pump to maintain a base pressure mbar.
-Ca ions in the ion trap originate from an integrated Ca target mounted near the trap.
-By ablating the target with 532 nm laser pulses, neutral Ca atoms are introduced to the ion trap where they are photoionized in a two-step process using 375 nm and 422 nm lasersPogorelov et al.
+Ca+ ions in the ion trap originate from an integrated 40Ca target mounted near the trap.
+By ablating the target with 532 nm laser pulses, neutral 40Ca atoms are introduced to the ion trap where they are photoionized in a two-step process using 375 nm and 422 nm lasersPogorelov et al.
 (2021).
-The generated Ca ions are then captured by the ion trap and Doppler-cooled by driving their 4S 4P cycling transition at 397 nm, as shown in the inset of Fig.
+The generated Ca+ ions are then captured by the ion trap and Doppler-cooled by driving their 42S1/2 42P1/2 cycling transition at 397 nm, as shown in the inset of Fig.
 2.
-Additional lasers at 866 nm and 854 nm are applied to repump the ions that decay to the 3D and 3D states, respectively, back to the cyclic cooling manifold.
-To the right of the main chamber, the molecular gas chamber is used to accumulate water vapor for the generation of CaOH molecular ions.
+Additional lasers at 866 nm and 854 nm are applied to repump the ions that decay to the 32D3/2 and 32D5/2 states, respectively, back to the cyclic cooling manifold.
+To the right of the main chamber, the molecular gas chamber is used to accumulate water vapor for the generation of CaOH+ molecular ions.
 When the gate valve is closed, this section is not pumped by the vacuum pumps attached to the main chamber and the water vapor pressure inside increases over time due to outgassing from the chamber wall.
@@
 III Molecule Generation and identification
-To generate CaOH for the photodissociation measurement, water vapor is introduced to the main chamber by opening the gate valve.
+To generate CaOH+ for the photodissociation measurement, water vapor is introduced to the main chamber by opening the gate valve.
 In our setup, the partial pressure of water vapor in the system cannot be directly measured.
 Instead, the increase in water vapor pressure in the main chamber is inferred from the change in total pressure estimated from the current of the ion pump.
-While waiting for the generation of CaOH, the gate valve is opened every 5 minutes for 1 second, which results in the total pressure estimated by the ion pump increasing from mbar to mbar.
+While waiting for the generation of CaOH+, the gate valve is opened every 5 minutes for 1 second, which results in the total pressure estimated by the ion pump increasing from mbar to mbar.
 As investigated by Okada et al.
-in 2003, the chemical reaction rate between HO and the trapped Ca ions increases when the ions are pumped into the 3D state by switching off the 866 nm repumping light during laser coolingOkada et al.
+in 2003, the chemical reaction rate between H2O and the trapped Ca+ ions increases when the ions are pumped into the 32D3/2 state by switching off the 866 nm repumping light during laser coolingOkada et al.
 (2003).
 Therefore, the 866 nm laser is periodically switched off for 1 second every 2 seconds in the molecule generation phase.
-The generation rate of one CaOH molecular ion is typically 10 to 15 minutes for the ion crystals of Ca ions investigated in this experiment.
-The above molecule generation process was found to produce predominantly CaOH by mass spectrometry using the method presented by Drewsen et al.
+The generation rate of one CaOH+ molecular ion is typically 10 to 15 minutes for the ion crystals of Ca+ ions investigated in this experiment.
+The above molecule generation process was found to produce predominantly CaOH+ by mass spectrometry using the method presented by Drewsen et al.
 Drewsen et al.
 (2004) (for details, refer to Appendix D).
-In this method, the unknown molecular mass can be inferred from the frequency of the axial center-of-mass (COM) motion of a 2-ion crystal of one Ca ion and one molecular ionKielpinski et al.
+In this method, the unknown molecular mass can be inferred from the frequency of the axial center-of-mass (COM) motion of a 2-ion crystal of one Ca+ ion and one molecular ionKielpinski et al.
 (2000); Drewsen et al.
 (2004) as
-where is the ratio between the atomic and molecular mass and is the COM motional frequency of a Ca ion crystal.
+where is the ratio between the atomic and molecular mass and is the COM motional frequency of a Ca+ ion crystal.
 These frequencies were measured by applying an oscillating voltage to one of the endcap electrodes.
@@
 3.
-All 65 measurements have masses consistent with CaOH, from which we conclude that the molecular ions generated in our procedure are predominantly this species.
+All 65 measurements have masses consistent with CaOH+, from which we conclude that the molecular ions generated in our procedure are predominantly this species.
 IV Photodissociation
-After verifying the generation of a single CaOH, the measurement of the photodissociation spectrum of CaOH is performed by applying femtosecond pulse trains produced by an OPA.
+After verifying the generation of a single CaOH+, the measurement of the photodissociation spectrum of CaOH+ is performed by applying femtosecond pulse trains produced by an OPA.
 The spatial beam profile and position of the OPA light are measured by a beam profiler mounted on a rail.
@@
 The radius of the OPA light at the ion crystal was also estimated by the beam profiler to be 150(10) µm.
-We assume that the dissociation of a CaOH molecule during illumination is memoryless due to the lack of intermediate electronic states.
+We assume that the dissociation of a CaOH+ molecule during illumination is memoryless due to the lack of intermediate electronic states.
 Thus, the dissociation probability after illumination time can be described by the cumulative distribution function (CDF) of an exponential distribution,
@@
 Here, is the photon energy, is the Planck’s constant, is the speed of light, and is the average intensity of the OPA light at the ion position with average power and beam radius .
-The experimental procedure is initialized by loading a one-dimensional ion chain with Ca ions into the ion trap for the generation of a single CaOH (see Fig.
-10).
-While Ca ions can be detected on an electron-multipling CCD (EMCCD) camera, the generated CaOH ions do not scatter light and are recognized as dark defects in the ion chain.
+The experimental procedure is initialized by loading a one-dimensional ion chain with Ca+ ions into the ion trap for the generation of a single CaOH+ (see Fig.
+11).
+While Ca+ ions can be detected on an electron-multipling CCD (EMCCD) camera, the generated CaOH+ ions do not scatter light and are recognized as dark defects in the ion chain.
 The number of dark ions is identified by comparing the ion crystal image with pre-calculated ion crystal configuration templates with different numbers of dark and bright ions (see Appendix C).
 After the appearance of a single molecular ion, the OPA light is applied to illuminate the molecular ion for up to 300 s.
-If the OPA light successfully recovers the original Ca from the CaOH ion, as shown in Fig.
+If the OPA light successfully recovers the original Ca+ from the CaOH+ ion, as shown in Fig.
 4, the illumination time for achieving photodissociation will be recorded.
@@
 based on Eq. (2) where denotes the total illumination time. Starting from a uniform prior distribution in the range of s, the posterior distribution after trials,
-reveals an estimation of . To ensure convergence of the posterior distribution, approximately 20 trials were performed at each wavelength.
+is normalized and reveals an estimation of . To ensure convergence of the posterior distribution, approximately 20 trials were performed at each wavelength.
 As shown in Eq.
@@
 V Two-photon dissociation
-In addition to the measurement of single-photon dissociation, we also investigated the photodissociation of CaOH at wavelengths around 500 nm.
+In addition to the measurement of single-photon dissociation, we also investigated the photodissociation of CaOH+ at wavelengths around 500 nm.
 To achieve a dissociation rate that is comparable to the single-photon measurements, the repetition rate of the OPA is increased to 10 kHz and the light intensity is increased by reducing the beam radius to 70(10) µm and increasing the average power to a few milliwatts.
@@
 Fitting the data to a power law function gives an exponent of , which is consistent with two-photon processes.
-Thus, the photodissociation of CaOH at wavelengths around 500 nm is assumed to be dominated by a two-photon dissociation channel via its unbound first electronic excited state.
+Thus, the photodissociation of CaOH+ at wavelengths around 500 nm is assumed to be dominated by a two-photon dissociation channel via its unbound first electronic excited state.
 The two-photon dissociation cross section is related to the dissociation rate by (see Appendix A.2)
@@
 6.(b).
-VI Discussion
-In summary, we measured the photodissociation spectrum of single trapped CaOH molecular ions sympathetically cooled in Ca ion crystals at nm and nm.
+VI Quantum-chemical modelling
+To interpret experimental measurements and predict theoretical spectroscopic parameters of CaOH+, we use state-of-the-art quantum-chemical methods.
+We employ wave-function approaches: the coupled cluster method restricted to single, double, and noniterative triple excitations, CCSD(T) Hampel, Peterson, and Werner (1992), to describe the ground electronic state and internally contracted multireference configuration interaction method restricted to single and double excitations, MRCISD Werner and Knowles (1988), to describe excited electronic states.
+Electronic orbitals are constructed using the augmented correlation-consistent polarized weighted core-valence quintuple- quality basis set (aug-cc-pwCV5Z-PP Hill and Peterson (2017) for Ca, aug-cc-pwCV5Z Peterson and Dunning (2002) for O, and aug-cc-pV5Z Dunning (1989) for H).
+We include the scalar relativistic effect in Ca using the small-core relativistic energy-consistent pseudopotential (ECP10MDF Lim, Stoll, and Schwerdtfeger (2006)) to replace the inner-shell electrons.
+Electronic structure calculations are performed with the Molpro Werner et al.
+(2012a, b) package of ab initio programs.
+Uncertainties of present energy calculations should not be larger than 1-2%Gronowski, Koza, and Tomza (2020).
+We optimize the ground-state geometry of linear CaOH+ and find the equilibrium internuclear distances bohr and bohr.
+Such an equilibrium geometry gives the rotational constant cm-1.
+Present theoretical harmonic frequencies for normal modes of CaOH+ are 3949.9 cm for OH stretching, 747.4 cm-1 for CaO stretching, and 481.2 cm-1 for doubly degenerate bending.
+The frequency for the OH mode is slightly larger as compared to our theoretical value of 3748 cm-1 and the experimental value of 3737.8 cm-1 Huber and Herzberg (1979) for a free OH.
+Additionally, using the finite-field approach, we predict the equilibrium permanent electric dipole moment of CaOH+ with respect to its center of mass to be 6.3 debye.
+This relatively large dipole moment originates from the partially charge-transferred nature of the ground-state CaOH+, which can be formally described as Ca2+OH-.
+Under room temperature (300K), the molecular ion is likely to be in its vibrational ground state (97% for CaO stretching).
+Using the CCSD(T) method, we predict that the dissociation energy of CaOH+ in the ground rovibrational level () into and OH () is 37807 cm-1, which corresponds to 264.5 nm.
+This value originates from 38763 cm-1 of the electronic contribution corrected by cm-1 of the zero point energy (ZPE) difference within the harmonic approximation, where the ZPE for CaOH+ and OH is 2830 cm-1 and 1874 cm-1, respectively.
+Figure 7 presents the ground and excited electronic states of the CaOH+ in the linear geometry as a function of the distance between Ca and the center of mass of OH.
+OH is described within the rigid-rotor approximation with the experimental internuclear distance of = 1.8332 bohr Drouin (2013).
+The one-photon dissociation of CaOH+ is driven by the allowed electric dipole transition from the ground state to the excited state, which is slightly repulsive around the ground-state equilibrium geometry and dissociates into Ca+ and OH.
+The corresponding vertical transition moment obtained with the MRCISD method is 0.8 debye.
+The photodissociation cross sections are given by the convolution of the transition moment function and the overlap of the ground rovibrational and excited continuum nuclear wave functions.
+Exact values of cross sections are sensitive to the precise location of the classical turning point of the state, which is, however, robustly predicted by the employed ab initio methods to be within the width of the ground rovibrational wave function.
+VII Discussion
+In summary, we measured the photodissociation spectrum of single trapped CaOH+ molecular ions sympathetically cooled in Ca+ ion crystals at nm and nm.
 The photodissociation processes at wavelengths in these two ranges are verified to originate predominantly from a single-photon and a two-photon dissociation channel, respectively.
@@
 Compared to the previous workOkada et al.
-(2006) on single-photon dissociation of CaOH by Okada et al.
-in 2006, we report a more precise measurement of the dissociation cross sections at various wavelengths which is consistent with the previous estimation of at nm.
-The measured single-photon dissociation spectrum is also consistent with the photodissociation of CaOH at 272 nm observed by Hansen et al.Hansen et al.
+(2006) on single-photon dissociation of CaOH+ by Okada et al.
+in 2006, we report a more precise measurement of the dissociation cross sections at different wavelengths which is consistent with the previous estimation of at nm.
+The measured single-photon dissociation spectrum is also consistent with the photodissociation of CaOH+ at 272 nm observed by Hansen et al.Hansen et al.
 (2012).
-As for the two-photon dissociation of CaOH around nm, the dissociation process is found to be significantly less efficient compared to the single-photon dissociation process, as expected since higher peak intensity is typically required for higher-order processes.
+As for the two-photon dissociation of CaOH+ around nm, the dissociation process is found to be significantly less efficient compared to the single-photon dissociation process, as expected since higher peak intensity is typically required for higher-order processes.
 Yet femtosecond lasers at such wavelengths have proven to be alternatives to continuous wave lasers with nm for photodissociation of the molecular ion.
 The question arises as to whether nanosecond lasers that are used for ablation loading in trapped atomic ion experiments at nmPogorelov et al.
-(2021) are able to dissociate CaOH.
-Achieving a mean photodissociation time of 1 minute for a single CaOH ion is estimated to require an average power of W, considering a laser with kHz, ns, and the light focused onto the ions with a beam waist of µm.
-The reported measurement is implemented on single CaOH molecular ions in small trapped ion crystals, which enables the relatively precise evaluation of the photodissociation cross section due to the repeatable positioning of the molecule.
+(2021) are able to dissociate CaOH+.
+Achieving a mean photodissociation time of 1 minute for a single CaOH+ ion is estimated to require an average power of W, considering a laser with kHz, ns, and the light focused onto the ions with a beam waist of µm.
+We employed advanced ab initio molecular electronic structure methods to characterize theoretically the ground-state properties of CaOH+ and to investigate the spectrum of excited electronic states involved in the measured photodissociation.
+We predicted the photodissociation threshold for CaOH+Ca++OH at 265 nm in very good agreement with the one-photon experimental observations.
+Efficient photodissociation is possible because of the coincidence of the slightly repulsive part of the excited electronic state close to its inner classical turning point, which is accessible by a decent electric dipole transition moment.
+The reported measurement is implemented on single CaOH+ molecular ions in small trapped ion crystals, which enables the relatively precise evaluation of the photodissociation cross section due to the repeatable positioning of the molecule.
 The scale of the system is also crucial for investigating a two-photon dissociation process that requires tightly focused light to achieve sufficient intensity.
-The reported photodissociation spectrum can be a basis for studying the rovibrational structure of CaOH via dissociation-based spectroscopy.
-Designed to investigate the interaction between femtosecond laser pulses and a linear ion crystal with a single CaOH, this experiment represents a precursor for performing quantum logic spectroscopy with polyatomic molecular ions Schindler (2019).
+The reported photodissociation spectrum can be a basis for studying the rovibrational structure of CaOH+ via dissociation-based spectroscopy.
+Designed to investigate the interaction between femtosecond laser pulses and a linear ion crystal with a single CaOH+, this experiment represents a precursor for performing quantum logic spectroscopy with polyatomic molecular ions Schindler (2019).
 Acknowledgements
-This research was funded by ERC-2020-STG 948893, ESQ Discovery project SDEF: State-dependent force spectroscopy for trapped ions, and FWF 1000 Ideas project TAI-798.
-The authors would like to acknowledge AQT for support with the ion trap and vacuum pumps, in particular Daniel Nigg and Georg Jacob; Markus Teller for assistance with vacuum work; electronics workshop technicians Kilian Prokop and Wolfgang Kuen for lab infrastructure support; Milan Oncak for discussions on quantum chemistry calculations; Christian Marciniak for general discussions; Marco Valentini for Ca lasers implementation; and the Quantum Optics & Spectroscopy Group and associated ion trapping groups at the Universität Innsbruck for general assistance.
+This research was funded by ERC-2020-STG 948893, ESQ Discovery project SDEF: State-dependent force spectroscopy for trapped ions, FWF 1000 Ideas project TAI-798, and the Foundation for Polish Science within the First Team programme.
+We acknowledge Poland’s high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational project PLG/2023/016115.
+The authors would like to acknowledge AQT for support with the ion trap and vacuum pumps, in particular Daniel Nigg and Georg Jacob; Markus Teller for assistance with vacuum work; electronics workshop technicians Kilian Prokop and Wolfgang Kuen for lab infrastructure support; Milan Oncak for discussions on quantum chemistry calculations; Christian Marciniak for general discussions; Marco Valentini for Ca+ lasers implementation; and the Quantum Optics & Spectroscopy Group and associated ion trapping groups at the Universität Innsbruck for general assistance.
 Author Contributions Z.W., S.W., B.F., and P.S.
@@
 performed the characterization of the OPA system.
-Z.W., S.W., and B.F.
+P.G.
+and M.T.
+performed electronic structure calculations.
+Z.W., S.W., B.F.
+and M.T.
 contributed to the manuscript.
-All authors contributed to the experimental setup and reviewed the manuscript.
+Z.W., S.W., V.P., M.I.M., E.M., G.M., R.N., B.F., and P.S.
+contributed to the experimental setup.
+All authors reviewed the manuscript.
 B.F.
@@
 Data and Code Availability
-The data and analysis code that support the findings of this study are openly available at https://github.com/quantum-molecules/dissociation_analysis.
+The data that support the findings of this study are openly available at http://doi.org/10.5281/zenodo.11109790.
 Appendix A Photodissociation model
@@
 We can therefore also estimate the mean dissociation time using the estimated dissociation probability for fixed times.
-These probabilities are then fitted to the exponential distribution CDF as indicated in Fig.
-7, e.g., for nm and nm.
+These probabilities are fitted to the exponential distribution CDF as indicated in Fig.
+8, e.g., for nm and nm.
 We estimate the photodissociation cross sections via this analysis for all data and compare them to the cumulative dissociation time, as shown in Fig.
-8.
+9.
 We find that the cross sections obtained from both methods agree within the uncertainties which indicates that the single- and two-photon dissociation processes are memoryless.
 Appendix B Experimental details
-B.1 Ca Lasers and trap parameters
+B.1 Ca+ Lasers and trap parameters
 During the photodissociation measurement, the cooling laser at 397 nm and the repumpers at 866 nm and 854 nm are constantly applied for Doppler cooling the ion crystal.
@@
 To verify this, the pulse duration and the spectrum of the OPA at nm were measured and the time-bandwidth product was inferred, as shown in Fig.
-9.
+10.
 To measure the pulse duration, we used an autocorrelator setup that splits the light from the OPA into two paths, where one path has an adjustable time delay.
@@
 Appendix C Ion crystal configuration identification
-The photodissociation experiment requires real-time evaluation of the ion crystal configuration such that the number of fluorescing (bright) Ca and non-fluorescing (dark) molecular ions can be monitored.
+The photodissociation experiment requires real-time evaluation of the ion crystal configuration such that the number of fluorescing (bright) Ca+ and non-fluorescing (dark) molecular ions can be monitored.
 Knowledge of the ion crystal configuration allows a script to run automatic measurements and take actions such as (re)loading the trap or emptying it if more than one dark ion is created.
-The ion crystal configuration is identified using only image data of the ions taken by an imaging system with magnification using an Andor iXon+ 885 camera with a pixel size of 8 µm.
+The ion crystal configuration is identified using solely image data of the ions taken by an imaging system with magnification using an Andor iXon+ 885 camera with a pixel size of 8 µm.
 For this process, a region of interest (ROI) in the image centered around the ions is converted into two normalized data sets: the column sum, where the ROI is summed along the columns of the image array; and the row sum, where the summation is done along the rows (see Fig.
-11).
+12).
 All analysis to identify the ion crystal configuration is based on these two data sets.
@@
 A threshold is then set to differentiate between an empty or a non-empty trap.
-The number of bright ions is counted from the row sum via a peak finder algorithm, which can determine the number and positions of bright ions without requiring a fluorescence-based threshold.
-This algorithm identifies the largest amplitude difference in a set of sorted local maxima (peaks) of row sum and uses this to classify peaks into two distinct groups: large-amplitude peaks originating from bright ions and small-amplitude peaks due to noise.
+The number of bright ions is counted from the column sum via a peak finder algorithm, which can determine the number and positions of bright ions without requiring a fluorescence-based threshold.
+This algorithm extracts and sorts local maxima of column sum over subsets of length comparable but smaller than the ion separation.
+From the largest amplitude difference and their amplitudes, the algorithm classifies peaks into two distinct groups: large-amplitude peaks originating from bright ions and small-amplitude peaks due to noise.
 Knowledge of the positions of the bright ion peaks can be utilized to identify the presence of dark ions for, e.g., determining the trap center position, as it requires an ion crystal of solely bright ions.
@@
 The desired ion configuration in a template for ions is given by the vector .
-The value of is assigned to for a bright Ca ion and for a dark molecular ion, for , e.g., the ion crystal configuration template which matches the row sum shown in Fig.
+The value of is assigned to for a bright Ca+ ion and for a dark molecular ion, for , e.g., the ion crystal configuration template which matches the row sum shown in Fig.
 12 is .
@@
 Here denotes the numerically calculated positionsJames (1998) of the ions with respect to the calibrated trap center position , is the waist of the Gaussian peaks (taken from the fit presented in Fig.
-11), and is an offset of the trap center position.
+12), and is an offset of the trap center position.
 The start and end positions, and , of the -th binning interval are chosen such that the binning and length of the template match those of the measured column sum data set.
 is the binned template of ions for the configuration .
-To find the particular ion crystal configuration, a large set of templates is created for all reasonable ion numbers and all possible corresponding ion configurations.
+To find the particular ion crystal configuration, a set of templates is created for all reasonable dark and bright ion numbers and all possible corresponding ion configurations.
 All templates are compared with the column sum data set using the comparison function
@@
 Appendix D Mass spectrometry
-A radio frequency (RF) “tickling” signal with an amplitude of 30 mV modulates the DC voltage on the endcap electrodes, which defines the axial confinement, and excites the COM motion of the ions which peaks when on resonance with the COM motional frequency.
+To determine the molecular mass, a radio frequency (RF) “tickling” signal with an amplitude of 30 mV modulates the DC voltage on the endcap electrodes which defines the axial confinement.
+The signal excites the COM motion of the ions which peaks when the signal frequency is on resonance with the COM motional frequency.
 Due to Doppler broadening, the ions fluoresce less when their motional state is excitedWineland, Drullinger, and Walls (1978).
@@
 In crystals with mixed ion species, the COM motional frequency depends not only on the different ion masses but also on the specific ion crystal configurationHome (2013).
-For instance, the two ion configurations (Ca, Ca, CaOH) and (Ca, CaOH, Ca) have a different COM motional frequency, with a deviation of 1.7 kHz in our system.
+For instance, the two ion configurations (Ca+, Ca+, CaOH+) and (Ca+, CaOH+, Ca+) have a different COM motional frequency, with a deviation of 1.7 kHz in our system.
 When applying an RF signal which is close to the COM motional frequency of the current ion configuration, the ions reconfigure their respective positions due to the induced motion.
@@
 Thus, we perform measurements with crystals consisting of exactly two trapped ions, as there exists only a single COM motional frequency due to symmetry.
-However, the dissociation measurements are performed with larger ion crystals to speed up the CaOH generation process.
-In order to prevent ion loss during the tickling scan, a Doppler cooling cycle of about is applied to the ions with the RF signal being turned off after each tickling measurement in the scan.
-For every single Ca and molecular ion pair, we performed and summed up to five measurement runs.
+However, the dissociation measurements are performed with larger ion crystals to speed up the CaOH+ generation process.
+To prevent ion loss during the tickling scan, a Doppler cooling cycle of about is applied to the ions with the RF signal being turned off after each tickling measurement in the scan.
+For every single Ca+ and molecular ion pair, we performed and summed up to five measurement runs.
 In order to determine the COM motional frequency from the images taken with the EMCCD camera, the measured photon counts at each pixel are summed in a small ROI around the ions.
 This ROI is wide along the y-axis of the images, where is the waist of a Gaussian fitted to the row sum dataset, as shown in Fig.
-11.
+12.
 Along the x-axis, this small ROI has a length which is 5 times that of a two ion crystal with an endcap electrode voltage of 205 V, i.e., 8.5 µm.
 The total photon counts integrated over the ROI as a function of the RF tickling frequency are normalized and the COM motional frequency is determined by fitting a Gaussian function to the data, as shown in Fig.
-12.
-The fitted COM motional frequencies from the 65 Ca and molecular ion pairs and the corresponding reference COM motional frequencies of Ca gave masses for the molecular ions ranging from amu.
-This confirms that all measured 65 molecular ions were CaOH.
-In order to estimate the statistical uncertainty of the measured COM motional frequency of trapped atomic Ca ions, we evaluated 34 data sets taken over the course of hours on two different days.
+13.
+The fitted COM motional frequencies from 65 Ca+ and molecular ion pairs and the corresponding reference COM motional frequencies of Ca+ gave masses for the molecular ions ranging from amu.
+This confirms that all measured 65 molecular ions were CaOH+.
+In order to estimate the statistical uncertainty of the measured COM motional frequency of trapped atomic Ca+ ions, we evaluated 34 data sets taken over the course of hours on two different days.
 For a voltage of 205 V applied to the endcap electrodes, a COM motional frequency of 526.8 kHz with a standard deviation of 200 Hz was determined.
-Since the errors in the fitted COM motional frequency of all 65 processed datasets of Ca and molecular ion pairs and all reference datasets are all smaller than 200 Hz, this value is utilized to obtain an upper limit for the uncertainty of COM motional frequencies.
+Since the errors in the fitted COM motional frequency of all 65 processed datasets of Ca+ and molecular ion pairs and all reference datasets are all smaller than 200 Hz, this value is utilized to obtain an upper limit for the uncertainty of COM motional frequencies.
 This results in an estimated uncertainty in the derived molecular ion masses of 0.1 amu.